# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v4

@@ Abstract
Recent minimalist modeling indicates that overdamped polar chiral active matter can support inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this article, we establish the statistical mechanical foundation for this emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as carrying non-dissipative phase rigidity supercurrents, a mapping we confirm empirically by demonstrating a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity. Generalizing this framework to three dimensions ($S^{1}\to S^{2}$), we show that polar alignment on the unit sphere is geometrically equivalent to the Gilbert damping term in spintronics, and that the two-dimensional Kuramoto coupling term naturally appears in the tangent-plane projection of spin relaxation. This constraint generates inertial spin waves (ferromagnetic magnons) from the overdamped active bath, recovering the macroscopic transport predicted by Toner-Tu theory without invoking microscopic inertia. Our results indicate that chiral active matter should strictly be considered a dissipative spintronic fluid, where phase gradient transport is ensured by the Goldstone modes of the underlying broken symmetry.
@@ I Introduction
A recent minimalist, two-dimensional chiral active matter model offers a solution to the paradox by empirically demonstrating that an overdamped chiral flock can drive a conservative inertial cascade that culminates in inviscid Euler turbulence, in the form of a stable, sloshing Onsager dipole in the dynamic agent ensemble, where topological defects interact long-range through the Kirchoff-Onsager hamiltonian [8] (Figure 1). The present article provides detailed theoretical explanation for the findings in Ref. [8].
As we shall demonstrate, the agent inertia that is engineered into several successful hydrodynamic models of chiral flocks [9, 10, 11] can be replaced by a natural synchronization stiffness that mimicks an inertial flow, clearly observable in our minimalist model.
@@ II Theoretical Foundation
In this letter, we shall consider Eq. (2) as the overdamped Langevin equation, following Ref. [19], which leads to the tilted washboard potential [17]. Given that our model’s agents are essentially massless, we shall derive a formal mapping to the resistively shunted junction model in the overdamped limit [20, 21].
@@ IV Reconciling the model with real, three-dimensional chiral flocks
Mathematically, the three-dimensional definitions preserved the symmetry group of the interaction while lifting the dimensionality, from which Kuramoto synchronization emerges without the ansatz.
@@ V Overdamped spin waves as ‘second sound’ in Toner-Tu theory
To fully capture the wave-like transport observed in Toner-Tu theories [9], we restore the inertial term neglected in the overdamped limit. We introduce a coefficient of turning inertia, or precession stiffness, $\gamma$:
@@
We note that this restoration of effective inertia is physically mandated by the conservation of angular momentum density: as the local order parameter $R$ vanishes, the collective moment of inertia disappears ($I_{\text{eff}}\propto R$), compelling the precession rate to diverge to sustain a finite interaction torque. This description remains valid only while the local timescale separation persists, breaking down precisely at the defect cores and at the boundaries ($R\rightarrow 0$) where the divergence signals the transition from a hydrodynamic superfluid to a strictly microscopic, overdamped gas.
To satisfy dimensional consistency ($\gamma]=L^{2}T^{-1}$) within this physical constraint, we define the inertial coefficient as:
$\gamma=\frac{a_{0}\xi^{2}}{R},$ (31)
which leads to a direct mapping to the sound speed of the topological gas, $c_{s}\approx v_{0}R$ [8],
$c_{\text{eff}}=v_{0}\sqrt{\frac{K_{\text{torque}}}{\gamma}}=v_{0}R,$ (32)
where $K_{\text{torque}}=a_{0}\xi^{2}R$.
Physically, the inverse dependency in Eq. (31) implies that at the flock boundary ($R\to 0$), as alluded to, the precession frequency diverges. The system transitions from a “heavy,” locked superfluid in the bulk to a massless, fast-spinning gas. This divergence creates the active bath required to sustain the topological protection of the inner Onsager dipole [8].
@@
By linearizing the polarized case (see Appendix A), the resulting spin waves are described, yielding the final total dispersion relation,
@@ VI.3 Ferromagnetic magnons
(2) the $k^{2}$ dependency in the dispersion relation corresponds to the Goldstone modes of the broken rotational symmetry (spin waves, $\omega(k)\propto k^{2}$). While a $k^{2}$ dependency can be modeled with odd viscosity [37], both the primary wave propagation for the Goldstone mode ‘second sound’ waves of Toner-Tu [9, 38] and phenomenological spin-wave models [11, 10] predict linear dispersion $\omega(k)\propto k$ (see Figure 6).
@@
The traditional hydrodynamic modes of transfer and the spintronic waves we propose to underpin information transfer are expected to coexist in a real chiral flock. Nevertheless, we expect regimes in which orientation (magnons) should persist when sound waves in density fail, and regimes where physical transport (agent velocity) poses clear limitations on the second sound propagation, but where the magnons are nevertheless able to transmit phase gradients at great speed. These cases offer ways to test the degree to which a chiral active matter is indistinguishable from a dissipative spintronic fluid.
The above notwithstanding, in our model, propagating magnons cause a high-frequency smearing that rapidly delocalizes information. As shown in Figure 6, the dispersive nature of the magnon mode acts as an early-warning system, allowing high-frequency ‘panic’ signals to propagate through the marginally synchronized bulk well in advance of the density wavefront [39]. However, the marginal synchronization is likely realized on the spintronic fluid’s dissipative boundary, while in the ordered bulk ($R\to 1$), the magnons are completely damped out. The spintronic chiral flock therefore does not turn or flow, but rather shatters and reforms, described in terms of shock merger in Ref. [8]. Here, the Josephson inequality (Eq. 25) defines a threshold between trapped (superconducting, $|\omega-\dot{\Psi}|<I_{c}$) and phase-slipping (dissipative, $|\omega-\dot{\Psi}|>I_{c}$) turbulence. This shattering behaviour at impact is directly supported in the literature [40], and observations of “anomalous diffusion” during turns [39] are highly consistent with the notion of agents turning too abruptly, then phase-slip $|\omega-\dot{\Psi}|>I_{c}$. In the spintronic fluid, turns are initiated at the boundaries of the ordered flock where it costs almost no energy to phase-slip. This chirality is provided by the active bath ($\Delta\omega$), directly supported by Ref. [41]’s finding that turning flocks nucleate at the tips. What is more, the natural Goldstone dynamics with ferromagnetic magnons that we uncover in our model is consistent with observations of starling flocks where phase gradients travel faster than the bulk sound speed [41], as well as the paradigm of scale-free correlations in starling flocks [42].
Given the close correspondence between a spintronic fluid description and the various hydrodynamic descriptions that has accumulated, the tilted washboard potential should remain valid for many chiral active matter systems. It, and the Goldstone dynamics it drives, follows directly from the geometry of alignment on the unit sphere, which, in the marginal synchronization regime, effectively forces the system to support such spin waves. Subsequently, aspects of chiral flocking, and the emergent Euler turbulence recovered in the model (Figure 1), are governed by the universality of the Landau-Lifshitz-Gilbert equation.
@@
Figure 6: Space-time plots of a viscous hydrodynamic mode (panel a) and a ferromagnetic magnon (panel b), generated by first decomposing an initial Gaussian ($\sigma=1)$ signal $u(x,0)$ into its wavenumber components $\hat{u}(k)$ using a fast Fourier transform (FFT). These components are then evolved in the frequency domain, with $\omega(k)=k-ik^{2}/2$ for the linear sound waves and Eq. (46) with $R=0.2$ and $a_{0}=0.5$ for the quadratic spin waves. Finally, the evolved signal is reconstructed via an inverse FFT, $u(x,t)=\mathcal{F}^{-1}[\hat{u}(k,t)]$, and the resulting amplitude is plotted as a heatmap to visualize the transport dynamics and relative damping of both modes.
VI.4 Toner-Tu theory
@@
With spintronic chiral active matter, we provide a theoretical foundation for spin wave-models of chiral flocking. Notably, Ref. [11], where the authors explicitly introduced a phenomenological inertial spin model for starling flocks. While the chiral flock is fundamentally dissipative, the turning dynamics effectively possess inertia [10, 38], which allows for the propagation of information as spin waves [37, 43]. This is the established, robust feature of polar chiral active matter [9, 44, 45]. By deriving the phenomenon directly from the agent interaction rules, via the Adler equation and the Landau-Lifshitz-Gilbert equation, we provide a natural microscopic foundation for the concept of elasticity in Toner-Tu field theory. This microscopic foundation allows us to isolate the Josephson critical current $I_{c}$ as the threshold (maximum allowed turning rate of an agent) that breaks the macroscopic rigidity. When $|\omega-\dot{\Psi}|>I_{c}$, the fluid locally melts into the active bath, at which point the ´second sound’ modes of Toner-Tu field theory likewise collapse. The inequality Eq. (25) may provide a quantitative threshold for the realization of flocking.
@@ VI.6 Disorder as source of order
Given the complete isomorphism between our minimalist model and disordered Josephson junction theory, our assertion that noise optimizes a system by pushing it towards criticality is supported by research into non-equilibrium quantum thermodynamics [25, 26], providing an avenue for spintronics to accurately model active matter.
@@
The key insight from our results lies in the origin of the description we have elucidated in this section. The 3D geometric constraints of polar, chiral agents lead directly to a two-dimensional Kuramoto coupling term in the tangent plane projection (Figure 4), and this action is entirely isomorphic to Gilbert damping in a magnetic field. The reason for this convergence is that both spintronics and chiral active matter obey the universality of the Landau-Lifshitz-Gilbert equation. Dry chiral active matter systens must, therefore, be strictly considered as dissipative spintronic fluids, where the active bath of the overdamped agents provides the thermodynamic pressure that stabilizes an inertial phase rigidity.
@@ VII Conclusion
By extending the model to three dimensions, we recovered the Landau-Lifshitz-Gilbert equation for an overdamped spintronic fluid. The information supercurrents that we have predicted and described in empirical terms transmit phase gradients across the chiral flock, sustaining Goldstone modes from the precession around the mean field of the 3D active bath. The resulting spin waves advect phase gradients, providing inertia to the inviscid Euler turbulence of the Onsager dipole. The “phantom inertia” discovered in Ref. [8], $\lambda=R^{2}$, is therefore physically identified as the precession stiffness of the order parameter in three dimensions.
If substantiated in future studies, the transition to active chiral turbulence must be considered the emergence of a topological quasiparticle gas, where macroscopic Onsager dipoles, in turn bound states of phase singularities, acquire effective inertia from the Goldstone modes (azimuthal precession) of the underlying broken symmetry.

### v5

@@ Abstract
Recent minimalist modeling indicates that overdamped polar chiral active matter can support inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this article, we establish the statistical mechanical foundation for this emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as carrying non-dissipative phase rigidity supercurrents, a mapping we confirm empirically by demonstrating a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity. Generalizing this framework to three dimensions ($S^{1}\to S^{2}$), we show that polar alignment on the unit sphere is geometrically equivalent to the Gilbert damping term in spintronics, and that the two-dimensional Kuramoto coupling term naturally appears in the tangent-plane projection of spin relaxation. This constraint generates inertial spin waves (ferromagnetic magnons) from the overdamped active bath, recovering the macroscopic transport predicted by Toner-Tu theory without invoking microscopic inertia. Our results show that all polar chiral agents who seek to align on the unit sphere strictly constitute a dissipative spintronic fluid, where phase gradient transport is ensured by the Goldstone modes of the underlying broken symmetry.
@@ I Introduction
A recent minimalist, two-dimensional chiral active matter model offers a solution to the paradox by empirically demonstrating that an overdamped chiral flock can drive a conservative inertial cascade that culminates in inviscid Euler turbulence, in the form of a stable, sloshing Onsager dipole in the dynamic agent ensemble, where topological defects interact long-range through the Kirchoff-Onsager hamiltonian [8] (Figure 1). However, in real chiral flocks, information and navigation is transmitted through second-sound waves [9], dispersive odd viscosity [10], or inertial spin waves [11], explanations that assume an inertial microscopic mass in the overdamped system. In the present article, we demonstrate that transport can instead be achieved by the natural synchronization stiffness in our model, which mimicks an inertial flow [8], offering an alternative to mass transport based on the flock’s orientation field.
@@ II Theoretical Foundation
In this article, we shall consider Eq. (2) as the overdamped Langevin equation, following Ref. [19], which leads to the tilted washboard potential [17]. Given that our model’s agents are essentially massless, we shall derive a formal mapping to the resistively shunted junction model in the overdamped limit [20, 21].
@@ IV Reconciling the model with real, three-dimensional chiral flocks
Mathematically, lifting the definition to three dimensions preserves the symmetry group of the interaction while expanding the phase space. From this higher-dimensional geometry, Kuramoto synchronization emerges naturally as a kinematic constraint that recovers the Lohe sphere model [36] through physical necessity.
@@ V Overdamped spin waves as ‘second sound’ in Toner-Tu theory
Having established the dissipative (even) term of the Landau-Lifshitz-Gilbert equation, we now derive the reactive (odd) component. This term arises from the conservation of the fluid’s intrinsic angular momentum density $\mathbf{J}$,
$\mathbf{J}=\lambda\mathbf{\Omega},$ (30)
where we identify the phase stiffness of the renormalized superfluid [8],
$\lambda=R^{2}.$ (31)
We then have,
$\frac{\partial\mathbf{J}}{\partial t}=\bm{\tau}_{\omega}+\bm{\tau}_{\text{G}},$ (32)
where $\bm{\tau}_{\omega}$ is the torque caused by intrinsic chirality $\omega_{i}$ and $\bm{\tau}_{\text{G}}$ is the Gilbert damping torque (Figure 4). For the inertial term, the precession rate $\partial\hat{n}/\partial t$, we have,
@@
where we used Eq. (31), and where we identify the inertial term of the Landau-Lifshitz-Gilbert equation,
$\gamma\equiv\frac{a_{0}\xi^{2}}{R}.$ (34)
Looking at the steps from Eqs. (30) to (33), we note that our model shares a spintronic superfluidic description with Volovik’s Helium 3He-A [37].
With Eq. (34), we express the Landau-Lifshitz-Gilbert equation,
[…]
where we immediately recognize the expressions for even (shear) and odd viscosity, $\nu_{E}$ and $\nu_{O}$,
$\nu_{E}=\Gamma=a_{0}R\xi^{2},\;\;\;\text{ and }\;\;\;\nu_{O}=\gamma=\frac{a_{0}\xi^{2}}{R}.$ (36)
The former represents Gilbert damping (Eq. 29) and the latter represents the parity (time-reversal symmetry) breaking Goldstone mode (Figure 5).
We map Eq. (35) to the standard Landau-Lifshitz-Gilbert equation,
[…]
where $\alpha$ is the dimensionless Gilbert damping parameter. Comparing this equation term-by-term with Eq. (35), we obtain the ratio of the damping term to the precession term, which isolates $\alpha$, via Eq. (36),
$\alpha=\frac{\Gamma}{\gamma}=\frac{a_{0}R\xi^{2}}{\left(\frac{a_{0}\xi^{2}}{R}\right)}=R^{2}.$ (38)
Thus, within the geometry of the model (Figure 4b), the Gilbert damping parameter $\alpha$ is equal to the order squared $R^{2}$. This allows us to express the viscosity terms as functions of the Gilbert damping term $\alpha$,
$\nu_{E}=\nu_{0}\alpha^{1/2},\;\;\;\text{ and }\;\;\;\nu_{O}=\nu_{0}\alpha^{-1/2},$ (39)
where $\nu_{0}\equiv a_{0}\xi^{2}$ is the bare interaction viscosity. We now note that the Gilbert damping term is dynamic and dependent on order squared, $\alpha=R^{2}$. Consequently, since a Kuramoto-coupling term is explicitly found as a projection of the Gilbert torque into the agent vector’s tangent plane, Eq. (39) represents the natural hydrodynamics of Kuramoto-synchronization [38].
@@
We observe that as the local order parameter $R$ vanishes, the shear viscosity disappears, compelling the precession rate to diverge to sustain a finite interaction torque. This description remains valid only while the local timescale separation persists, breaking down precisely at the defect cores and at the boundaries ($R\rightarrow 0$) where the divergence signals the transition from a hydrodynamic superfluid to a strictly microscopic, overdamped gas. To describe this aspect of the model, we linearize the polarized agent (see Appendix A), yielding the final total dispersion relation,
@@ VI.3 Ferromagnetic magnons
(2) the $k^{2}$ dependency in the dispersion relation corresponds to the Goldstone modes of the broken rotational symmetry (spin waves, $\omega(k)\propto k^{2}$), driving ferromagnetic magnons.
Here, we note that spin waves with $k^{2}$ dependency can be modeled with odd viscosity [10], and our model provides such odd viscosity (Eq. 39), only that it is purely emergent from flock coherence rather than a material quality, and it is only realized in the marginal synchronization regime.
@@
Moreoever, both the primary wave propagation for the Goldstone mode ‘second sound’ waves of Toner-Tu [9, 40] and phenomenological spin-wave models [11, 41] predict linear dispersion $\omega(k)\propto k$ (see Figure 6). This mode, and the foregoing, are expected to coexist with the spintronic waves that we propose underpins information transfer in real chiral flocks. Nevertheless, we expect regimes in which orientation (magnons) should persist when sound waves in agent density fail, and regimes where agent velocity poses limitations on the resulting density transport, but where the magnons are nevertheless able to transmit phase gradients at great speed.
@@
Figure 6: Space-time plots of a viscous hydrodynamic mode (panel a) and the ferromagnetic magnon predicted by the present article (panel b), generated by first decomposing an initial Gaussian ($\sigma=1)$ signal $u(x,0)$ into its wavenumber components $\hat{u}(k)$ using a fast Fourier transform (FFT). These components are then evolved in the frequency domain, with $\omega(k)=k-ik^{2}/2$ for the linear sound waves and Eq. (53) with $R=0.2$ and $a_{0}=0.5$ for the quadratic spin waves. Finally, the evolved signal is reconstructed via an inverse FFT, $u(x,t)=\mathcal{F}^{-1}[\hat{u}(k,t)]$, and the resulting amplitude is plotted as a heatmap to visualize the transport dynamics and relative damping of both modes.
The above notwithstanding, in our model, propagating magnons cause a high-frequency smearing that rapidly delocalizes information. As shown in the theoretical predictions in Figure 6, the dispersive nature of the magnon mode acts as an early-warning system, allowing high-frequency ‘panic’ signals to propagate through the marginally synchronized bulk well in advance of the density wavefront [42]. However, the marginal synchronization is likely realized on the spintronic fluid’s dissipative boundary, while in the ordered bulk ($R\to 1$), the magnons are completely damped out. The spintronic chiral flock therefore does not turn or flow, but rather shatters and reforms, described in terms of shock merger in Ref. [8]. Here, the Josephson inequality (Eq. 25) defines a threshold between trapped (superconducting, $|\omega-\dot{\Psi}|<I_{c}$) and phase-slipping (dissipative, $|\omega-\dot{\Psi}|>I_{c}$) turbulence. This shattering behaviour at impact is directly supported in the literature [43], and observations of “anomalous diffusion” during turns [42] are highly consistent with the notion of agents turning too abruptly, then phase-slip $|\omega-\dot{\Psi}|>I_{c}$. In the spintronic fluid, turns are initiated at the boundaries of the ordered flock where it costs almost no energy to phase-slip. This chirality is provided by the active bath ($\Delta\omega$), directly supported by Ref. [44]’s finding that turning flocks nucleate at the tips. What is more, the natural Goldstone dynamics with ferromagnetic magnons that we uncover in our model is consistent with observations of starling flocks where phase gradients travel faster than the bulk sound speed [44], as well as the paradigm of scale-free correlations in starling flocks [45].
Given the close correspondence between a spintronic fluid description and the various hydrodynamic descriptions that has accumulated, the tilted washboard potential should remain valid for many chiral active matter systems. It, and the Goldstone dynamics it drives, follows directly from the geometry of alignment on the unit sphere, which, in the marginal synchronization regime, effectively forces the system to support such spin waves. Subsequently, orientation transport in chiral flocks, including the inviscid Euler flow in the azimuthal turning inertia (Figure 1), are governed by the universality of the Landau-Lifshitz-Gilbert equation.
VI.4 Toner-Tu field theory
@@
With spintronic chiral active matter, we provide an alternative origin and theoretical foundation for spin wave-models of chiral flocking. Notably, Ref. [11], where the authors explicitly introduced a phenomenological inertial spin model for starling flocks. While the chiral flock is fundamentally dissipative, the turning dynamics effectively possess inertia [41, 40], which allows for the propagation of information as spin waves [10, 46]. This is the established, robust feature of polar chiral active matter [9, 47, 48]. By deriving the phenomenon directly from the agent interaction rules, via the Adler equation and the Landau-Lifshitz-Gilbert equation, we provide a natural microscopic foundation for the concept of elasticity in Toner-Tu field theory. This microscopic foundation allows us to isolate the Josephson critical current $I_{c}$ as the threshold (maximum allowed turning rate of an agent) that breaks the macroscopic rigidity. When $|\omega-\dot{\Psi}|>I_{c}$, the fluid locally melts into the active bath, at which point the ´second sound’ modes of Toner-Tu field theory likewise collapse. The inequality Eq. (25) may provide a quantitative threshold for the realization of flocking.
@@ VI.6 Disorder as source of order
Given the complete isomorphism between our minimalist model and disordered Josephson junction theory, we assert that noise optimizes a system by pushing it towards criticality (azimuthal precession in the marginal synchronization regime), a notion that is supported by research into non-equilibrium quantum thermodynamics [25, 26], found in oscillatory systems that likewise adhere to the universality of the Landau-Lifshitz-Gilbert equation.
@@
The key insight from our results lies in the origin of the aforementioned. The 3D geometric constraints of polar, chiral agents lead directly to a two-dimensional Kuramoto coupling term in the tangent plane projection (Figure 4), and this action is entirely isomorphic to Gilbert damping in a magnetic field. The reason for this convergence is that both spintronics and chiral active matter obey the universality of the Landau-Lifshitz-Gilbert equation. Dry chiral active matter systems must, therefore, be strictly considered as dissipative spintronic fluids, where the active bath of the overdamped agents provides the thermodynamic pressure that stabilizes an inertial phase rigidity.
@@ VII Conclusion
By extending the model to three dimensions, we recovered the Landau-Lifshitz-Gilbert equation for an overdamped spintronic fluid. The information supercurrents that we have predicted and described in empirical terms therefore transmit rigidity across the chiral flock, sustaining Goldstone modes from the precession around the mean field of the 3D active bath. The resulting spin waves advect phase gradients, providing inertia to the inviscid Euler turbulence of the Onsager dipole. The “phantom inertia” discovered in Ref. [8], $\lambda=R^{2}$, is therefore physically identified as the precession stiffness of the order parameter in three dimensions.
If substantiated in future studies, the transition to active, dry chiral turbulence must be considered the emergence of a topological quasiparticle gas, where macroscopic Onsager dipoles, in turn bound states of phase singularities, acquire effective inertia from the Goldstone modes (azimuthal precession) of the underlying broken symmetry.